OpenStudy (aroub):

Given a triangle ABC that is right-angled at point A .The hypotenuse BC is fixed while A varies. Find the locus of point A.

hartnn (hartnn):

any ideas ?

hartnn (hartnn):

after solving u get a CIRCLE

OpenStudy (aroub):

No, this one no idea at all!! Lol, I was about to say not circle.. Umm, how? They didn't say at a given distance or a given point, right?

hartnn (hartnn):

and we can use exactly the same diagram we used in last question

hartnn (hartnn):

here, just the condition is different

OpenStudy (aroub):

Yeah, but don't use it please!

hartnn (hartnn):

gues which theorem we gonna use when its given 'a right triangle' ?

hartnn (hartnn):

*guess

OpenStudy (aroub):

I always think of Pythagorean when it comes to right triangles

hartnn (hartnn):

that is absolutely correct!

OpenStudy (aroub):

But what does this theorem got to do with the locus?

hartnn (hartnn):

using pythagoras theorem and distance formula , we get the equation of locus in the form , \(x^2+y^2=a^2\) which is the equation of circle.

OpenStudy (aroub):

please please don't use them here.. because they're more confusing! And I never took them. Do you have any other way? If not.. Just leave this question aside =)

hartnn (hartnn):

* \(\\~ \large (x-h)^2+(y-k)^2=a^2\)

hartnn (hartnn):

question left ...

OpenStudy (aroub):

In the end it's circle right?

hartnn (hartnn):

absolutely!